OpenStudy (anonymous):

A map has a scale of 1 in : 25 mi. Two cities are 175 miles apart. How far apart are they on the map?

OpenStudy (anonymous):

\[\left( i \right)/\left( 25m \right) \times \left( 175m \right) \]

Directrix (directrix):

1inch is to 25 miles as d inches is to 175 miles 1/25 = d/ 175 d = 7 inches 7 inches on the map represents 175 miles